OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

@jigglypuff314 do you understand this?

jigglypuff314 (jigglypuff314):

a perpendicular line has a flipped opposite slope so first figure out what is the slope of the line given and then take the flipped opposite of it like if given slope = 3 then perpendicular slope = -1/3

OpenStudy (anonymous):

So, how would I figure that? Solve for x? and get Y = x?

jigglypuff314 (jigglypuff314):

it's already in slope intercept form y = mx + b where m = slope :)

OpenStudy (anonymous):

OH! Okay! XD Wow.

OpenStudy (anonymous):

\[Y = 2x + 6\] \[Y = -\frac{ 1 }{ 2 } + 3\] \[Y = -\frac{ 1 }{ 2 } + 6\] \[Y = 2x + 3\]

OpenStudy (anonymous):

I think A, but I'm confused. Functions are strange..

jigglypuff314 (jigglypuff314):

well, what did you get as the slope for y = 2x +2 ? :)

OpenStudy (anonymous):

2x?

jigglypuff314 (jigglypuff314):

close :) It would be just 2 now, what would be the flipped opposite of 2?

OpenStudy (anonymous):

-2

jigglypuff314 (jigglypuff314):

don't forget flipped so -1/2 :) now, can you make an equation out of slope = -1/2 and point (6, 3) ?

OpenStudy (anonymous):

I was about to say that! and its C! :D

jigglypuff314 (jigglypuff314):

Correct! :)

OpenStudy (anonymous):

Thank you so much!

jigglypuff314 (jigglypuff314):

You're Welcome ^_^